OpenStudy (anonymous):

Factoring: (a+b)^2 - (a-b)^2 I got (a+b-a-b)(a+b+a-b) but... then what? (0)(2a)=0???

hartnn (hartnn):

the first factor needs revision : it should be (a+b-a+b) check it....

OpenStudy (anonymous):

well i would first write it out like this: \[(a+b)(a+b)-(a-b)(a-b)\] do you see why that's correct instead of what you first wrote?

OpenStudy (anonymous):

(0)(2a)=0 is true

OpenStudy (anonymous):

@hartnn I don't understand that... its (A-B)(A+B), right? So if A is (a+b) and B is (a-b) , the first factor should be (a+b-a-b) . If that's wrong, could you explain that to me? @agentc0re Could you explain that?

OpenStudy (anonymous):

@85295james can you tell me why?

hartnn (hartnn):

A-B = (a+b)-(a-b) = (a+b -a- (-b) ) = (a+b-a+b)

OpenStudy (anonymous):

Ah. Thank you, that makes more sense...

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

0 X anyting is O lol

OpenStudy (anonymous):

i know:)

hartnn (hartnn):

so did u finally get 4ab as your answer ??

OpenStudy (anonymous):

Yep. Thanks everyone!

OpenStudy (anonymous):

your welcome